### v1

# Neural Networks and Friction: Slide, Hold, Learn

Abstract
In this study, it is demonstrated that Recurrent Neural Networks (RNNs), specifically those utilizing Gated Recurrent Unit (GRU) architecture, possess the capability to learn the complex dynamics of rate-and-state friction laws from synthetic data. The data employed for training the network is generated through the application of traditional rate-and-state friction equations coupled with the aging law for state evolution. A novel aspect of our approach is the formulation of a loss function that explicitly accounts for initial conditions, the direct effect, and the evolution of state variables during training. It is found that the RNN, with its GRU architecture, effectively learns to predict changes in the friction coefficient resulting from velocity jumps, thereby showcasing the potential of machine learning models in understanding and simulating the physics of frictional processes. The implications of these findings for the field of geophysics and fault mechanics are discussed, highlighting the model’s proficiency in capturing the intricate behavior of seismic slip phenomena.
friction, rate-and-state-dependent friction, neural networks, recurrent neural networks, GRU
preprint: AAPM/123-QED
I Introduction
Understanding the dynamics of friction is paramount across several disciplines, including mechanical engineering, civil engineering, and geophysics [1, 2]. In mechanical and civil engineering, precise knowledge of frictional forces is crucial for the design and analysis of machines, structures, and materials under varying operational conditions [3]. In geophysics, friction dynamics play a critical role in earthquake mechanics, influencing the initiation, propagation, and arrest of seismic slips along faults [4].
Friction modeling [5, 6] aims to capture essential features observed in experimental settings [7], notably the direct effect and healing [2]. The direct effect, also known as the velocity strengthening behavior, refers to the immediate increase in frictional resistance with an increase in sliding velocity. This phenomenon can be described by the empirical equation [4]:
$\mu=\mu_{0}+a\log\left({V\over V_{ref}}\right)\,,$ (1)
where $\mu$ is the friction coefficient, $\mu_{0}$ is the friction coefficient at a reference velocity $V_{ref}$, $V$ is the sliding velocity, and $a$ is a positive constant representing the direct effect’s sensitivity to velocity changes [8].
Healing, also referred to as state evolution effect, illustrates the time-dependent increase in frictional resistance during stationary contact, attributed to the contact area’s growth at asperity contacts or chemical bonding at the interface [2]. This process is often modeled by the evolution of an internal variable $\theta$ (termed “state”), with the simplest form being:
$\dot{\theta}=1-\frac{\theta V}{D_{c}}\,,$ (2)
where $D_{c}$ is a critical slip distance over which the state variable evolves significantly [9]. The state is interpreted as a measure of effective age of contacts between the asperities of the opposed surfaces [8], and it therefore has units of time.
Numerous models have been proposed to describe frictional behavior, from classical laws like Amontons-Coulomb friction [2], which simplifies friction as a constant proportion of the normal load, to more complex models incorporating velocity and time-dependent effects. However, these models often fall short in capturing the full range of observed behaviors [4, 10].
Rate-and-state friction (RSF) laws stand out by explicitly addressing the direct effect and healing, thereby providing a more comprehensive framework for understanding the seismic cycle [11]. These laws are formulated as:
$\mu=\mu_{0}+a\ln\left(\frac{V}{V_{ref}}\right)+b\ln\left(\frac{V_{ref}\theta}{D_{c}}\right)\,,$ (3)
where $b$ is a constant that, along with $a$, determines the frictional response to changes in sliding velocity and state evolution [12]. Despite their advantages, RSF laws are not without limitations. Identifying the parameters $a$, $b$, and $D_{c}$ from experimental data poses significant challenges due to the complex interplay between different physical processes at the frictional interface [4].
Criticism has been levied against the use of the ”state” variable in RSF laws for its empirical nature and lack of a clear physical basis, leading to calls for models that include more explicit physical mechanisms [13, 14]. To begin, the parameters that appear in eq. 3 are assumed to be constants, and they may not [15, 16, 17, 18]. Furthermore, the need to regularize these laws to address numerical instabilities and ensure the physical realism of simulations has been highlighted [19, 20]. Recently, laboratory [21] and virtual experiments [22, 23] have shown the limitations of RSF laws even in highly controlled interface conditions.
Figure 1: Artistic representation of the magnified rough interface between two plates in sliding contact, where asperities grinding give rise to debris particle formation.
Given these challenges and the limitations of existing models, there is a clear need to explore new methodologies for understanding friction dynamics. This necessity is particularly acute for modeling phenomena that span multiple scales of time and space [24]. Herein lies the potential of leveraging advanced computational approaches, such as Recurrent Neural Networks (RNNs) [25], to offer new insights and overcome the limitations of traditional friction laws.
Neural networks have revolutionized the field of machine learning, offering powerful tools for modeling complex, non-linear relationships in data across a myriad of disciplines [26]. At their core, neural networks are composed of layers of interconnected nodes or ”neurons,” each capable of performing simple computations. Through the process of training, these networks adjust their internal parameters to minimize the difference between their output and the desired outcome, effectively learning to map inputs to outputs [27].
Among the various architectures of neural networks, Recurrent Neural Networks (RNNs) stand out for their ability to process sequential data. Unlike feedforward neural networks, RNNs possess a form of memory that allows them to incorporate information from previous inputs into the current processing step [28]. This characteristic makes RNNs particularly suited for tasks involving time series data or any context where the sequence of data points is crucial.
Two notable advancements in RNN architecture are the Gated Recurrent Unit (GRU) and Long Short-Term Memory (LSTM) networks. Both GRU and LSTM address a fundamental limitation of basic RNNs known as the vanishing gradient problem, which makes it difficult for RNNs to learn dependencies between events that occur at long intervals in the input data [27]. GRUs simplify the LSTM architecture while retaining its ability to capture long-term dependencies, making it both efficient and powerful for a wide range of tasks [28].
The application of RNNs, GRUs, and LSTMs has been widespread and transformative across many fields, including physics [29]. Particularly, their success in modeling path-dependent plasticity of materials highlights their potential [30, 31, 32, 33, 34]. Path-dependent plasticity is a critical aspect in the study of materials science, where the history of material deformation affects its current and future state. RNNs have been shown to effectively model these complex relationships, capturing the history-dependent nature of material behavior [35, 36].
This success in modeling path-dependent phenomena in materials science provides a strong foundation for the belief that similar methodologies can be applied with success to the study of friction. Friction, much like material plasticity, is inherently dependent on the history of contact and slip between surfaces, characterized by complex state-dependent behaviors such as the direct effect and healing. The capability of RNNs to capture and learn from sequential, history-dependent data makes them promising candidates for advancing our understanding and modeling of friction dynamics. Their ability to learn from data where previous states influence future behavior aligns well with the challenges presented by rate-and-state friction laws, suggesting that these neural network models could offer new insights and predictive capabilities beyond traditional friction models.
The goal of this short communication is to show that RNNs, based on GRUs, can perform at least as well as RSF laws, the aging law in particular. This approach is essentially different from the one used recently by Ishiyama and collaborators [37], who used machine learning (random forest [38]) to improve the estimation of RSF parameters eq. 3: machine learning is not deployed to improve an empirical approximate law, but to replace it altogether, down the road, in larger-scale simulations[14, 39, 40]. It also differs from the use made by Wang et al [41] and Karimpouli et al [42], who used machine learning techniques to devise impending rupture predictions based on acoustic emissions recorded in lab settings. Herein, we assume a virtual interface that is perfectly modeled by a traditional slip-law eqs. 3 and 2, for which data for training and verification can be generated easily. We use those data to train a custom GRU network that learns the dynamics without internal variables, since the state is not a physical quantity that can be measured experimentally (assuming it even exists as such). In reality, the complexities of interface physics [43] are such that probably no simplified internal variable can be defined as long as any real one aims at capturing a realistic degree of complexity We presume implicitly that sliding amount and sliding rate are the sole driving forces of the phenomenon, and whatever internal variables there may be can be internally codified in the parameters of the net. The figure of merit will not be the dynamic friction coefficient itself, but variations in the friction coefficient due to sliding. This decision is motivated the practical inability to surmise the time evolution of the friction coefficient except if proper initial conditions are supplied.
The structure of the paper is as follows: the methods section (1) introduces the GRU architecture, (2) describes the generation of synthetic data for training and verification, and (3) explains the choice of training hyperparameters, including the physics-informed loss function. The next section describes the outcome of the training process, which details performance evaluation, and then assesses the model’s predictive accuracy on independent datasets. The discussion section analyzes the findings and their implications. The final remarks section summarizes the study’s contributions and suggests future research directions.
II Methods
II.1 Neural networks for sequences: gated recurrent units
RNNs are engineered to handle sequential data thanks to their ability to maintain a memory of previous inputs by feeding the output of a neuron back onto itself [25]. This recursive nature allows making predictions based on the information accumulated from prior inputs. However, standard RNNs often struggle with long-term dependencies due to the vanishing gradient problem, where gradients become too small or too large for the network to learn effectively over many time steps.
Gated Recurrent Units (GRUs) are an evolution of the traditional RNN architecture, designed to mitigate the vanishing gradient problem and improve the network’s ability to learn from long sequences. GRUs introduce gating units that control the flow of information. These gates “decide” which information is relevant to keep or discard as the sequence progresses, enabling the network to maintain a longer memory and capture dependencies over longer sequences.
The GRU mechanism is articulated mathematically through [28]:
  • Update gates:
    $z_{t}=\sigma(W_{z}x_{t}+U_{z}h_{t-1}+b_{z})\,.$ (4)
  • Reset gates:
    $r_{t}=\sigma(W_{r}x_{t}+U_{r}h_{t-1}+b_{r})\,.$ (5)
  • Candidate states:
    $\tilde{h}_{t}=\tanh(Wx_{t}+U(r_{t}\odot h{t-1})+b)\,.$ (6)
  • Final state updates:
    $h_{t}=(1-z_{t})\odot h_{t-1}+z_{t}\odot\tilde{h}_{t}\,.$ (7)
In these equations, $x_{t}$ represents the input at time step $t$; $h_{t}$ is the hidden state at time $t$; $\sigma$ denotes the sigmoid activation function (to ensure the gate activations are constrained between 0 and 1); $\tanh$ is the hyperbolic tangent function (introducing non-linearity); $W$, $U$, and $b$ are the weights and biases associated with the input and hidden states, respectively; and $\odot$ signifies element-wise multiplication.
By merging the forget and input gates into a single update gate, GRUs simplify the LSTM architecture, reducing the model’s complexity and computational demands while maintaining robust performance.
II.2 Synthetic data generation
As mentioned earlier, we do not consider the state $\theta$ to be a necessary variable, thus it will not be used henceforth.
To generate the dataset for training and verifying our model, we start by defining a traditional RSF model with specific parameter values, grounded in the fundamental equations presented earlier (eq. 3), with parameter values set as follows:
  • Friction coefficient, $\mu_{0}=0.5$
  • Rate-state parameter, $b=0.015$ (unitless)
  • Direct effect parameter, $a=0.005$ (unitless)
  • Reference velocity, $V_{\text{ref}}=1.0\times 10^{-4}$ m/s
  • Characteristic slip distance, $D_{c}=0.2\times 10^{-3}$ m
Using these values, we simulate 15000 random sequences to reflect diverse frictional sliding scenarios. Each sequence is designed to simulate a total sliding distance of $30D_{c}$ and is completed within 0.013 seconds. The number of velocity jumps within each sequence is randomly determined, drawing from a uniform distribution with values 3, 4, 5. It is assumed that each sequence initiates from a non-zero velocity, with the initial state of the interface corresponding to the steady state for that velocity, as defined by the RSF laws: $\theta_{ss}=D_{c}/V_{0}$.
Velocity jumps within these sequences occur at random intervals, determined by sampling times from a uniform distribution across the interval 0, 1, with 0 marking the start and 1 the end of the sliding protocol. Each sequence incorporates a hold period where the velocity drops to zero, simulating intervals where sliding stops until the subsequent jump.
To compute the velocities between jumps, we divide the total sliding distance by the time spans between these jumps, ensuring that the velocities remain constant in these intervals. Thus, we define the “velocity protocol” for each sequence, i.e., virtual experiment.
Having the velocity protocol, we proceed to calculate the evolution of the state variable $\theta$ using eq. 2 and initial condtions $V(t=0)=V_{0}$ and $\theta(t=0)=\theta_{ss}$. With the velocity protocol and the evolving state variable at hand, we can then compute the friction coefficient’s evolution over time, as described by the RSF model (eq. 3).
The dataset comprises pairs of “features” and “targets”. Features represent the velocity protocol sampled at 100 equally-spaced instants during the duration of each virtual experiment, while targets correspond to variations of the friction coefficient sequence at these same instants. What is the reason for working with variations and not with the actual values of friction coefficient? There is no way to predict the initial values of friction based on these datasets, they have to be provided as initial conditions over which the variations are superimposed.
For training, we normalize using the following characteristic values: $\Delta\mu_{ch}=0.01$ for the friction coefficient change and $V_{ch}=0.3$ m/s for the velocities. $20\%$ of the sequences are put aside to test the network once it is trained.
II.3 Network details
Mathematica’s neural networks package has been used.
II.3.1 Architecture
The main net, which is aimed at replacing eq. 3 and whose parameters must be learned via training, is made up by a GatedRecurrentLayer, followed by a folded LinearLayer to transforms the output of the GRU through a linear transformation. Despite its simplicity, this network combines memory and transformation capabilities, make it well-suited to our goal of understanding and predicting the dynamic behavior of rate-and-state friction laws from sequential data.
Figure 2: Scheme of main part of the network
II.3.2 Training hyperparameters
We employ the ADAM optimization method due to its adaptive learning rate capabilities, which helps in converging to the optimal solution more quickly and efficiently than conventional gradient descent methods. A learning rate of 0.001 is chosen as a starting point because it strikes a balance between fast convergence and the risk of overshooting the minimum of the loss function. This value is generally considered a good default and has been proven effective across a wide range of tasks. In order to ensure that each training batch has enough data to represent the dataset’s diversity while keeping the computational demand manageable, we choose to use a batch size of 32. This size is large enough to benefit from computational optimizations during training yet small enough to avoid the potential pitfalls of training with very large batches, such as poor generalization. Finally, we use 50 epochs to allow the model opportunity to learn from the data without excessive training time. This limit helps prevent overfitting by avoiding too many epochs over the training dataset, promoting a model that generalizes well to unseen data.
II.3.3 Loss function
Inspired by physics-informed neural networks, we devise a loss function to guide the optimization of the parameters of the network. The conditions that observations impose over the friction coefficient are the direct effect, i.e., logarithmic dependence of velocity magnitude, and logarithmic healing, i.e., logarithmic increase of the static friction coefficient with time (during holds)
${\partial\mu\over\partial\log V}=\mathrm{c}_{1}\,,\quad{\partial\mu\over\partial\log t}=\mathrm{c}_{2}\,,$ (8)
where $\mu_{s}$ is the static friction coefficient, $c_{1}$ and $c_{2}$ are interface-dependent constants. The latter correspond to $a$ and $(a-b)$ (eq. 3), but we are assuming we do not know that the interface is modeled by these parameters, so shall impose that the change of these relative variations is constant, not that the variations themselves are of a given magnitude. There seems to be a real possibility [18] that these constants may change over time, so, to adapt the framework to this fact, we chose not to enforce the same constants over all the protocols but individually for each sequence, by not mixing information from other different series.
Apart from these two conditions, we must impose that there are no variations until the first jump (as we took for granted that each experiment starts from steady-state conditions) and that the average error is small, in order for the prediction to follow also the general trend in the target data.
Let $\delta\bm{\mu}$ and $\delta\bm{\hat{\mu}}$ to represent the vector of normalized predictions and targets (respectively) for a given vector of features (velocity protocol) $\bm{V}$, i.e., $\delta\bm{\mu}=\mathsf{NN}(\bm{V}/V_{ch})$. See that $\delta\bm{\mu},\,\delta\bm{\hat{\mu}},\,\bm{V}\in\mathbb{R}^{N}$ (and $N=100$ in this case). The first part of the loss is the mean-squared error:
$\mathsf{loss}_{1}={||\delta\bm{\mu}-\delta\bm{\hat{\mu}}||_{2}^{2}\over N}\,,$ (9)
where |||||$\cdot$||| | ⋅ | | represents the $L_{2}$ norm of the vector and $N=100$ is the number of elements in the sequence. In overall graph for training, fig. 6, this operation is implemented by the neuron labeled $\mathsf{MS}$.
The second part of the loss represents the difference between the initial values of predictions and targets,
$\mathsf{loss}_{2}=\left(\delta\mu_{1}-\delta\hat{\mu}_{1}\right)^{2}\,,$ (10)
where the sub-index $1$ refers to the first element of every sequence. The implementation of this operation is depicted in fig. 3: it takes the first sequence value of both targets (Targets) and predictions (Predictions), subtract each other and square the difference. Thus, the output of the net is eq. 10.
Figure 3: Scheme of sub-net that implements comparison of initial condition.
The third term of the loss is meant to enforce the direct effect: a sudden peak in friction induced by a jump in leading velocity (which, moreover, is known to be proportional to the logarithm of the ratio of velocities before and after the jump, see section and I and appendix A). Assume that there are $N_{J}$ jumps in a given protocol, the term that measures the error between the one predicted by the network and the real one is
$\mathsf{loss}_{3}=\sum_{i=1}^{N_{J}}\left\{\Delta(\delta\mu)_{J_{i}}-\Delta(\delta\hat{\mu})_{J_{i}}\right\}^{2}\,,$ (11)
the symbol $i$ indexes each jump, while $J_{i}$ refers to the entry of the vector where the jump begins, hence the meaning of the notation $\Delta(\delta\mu)_{J_{i}}=\delta\mu_{J_{i}+1}-\delta\mu_{J_{i}}$. The network depicted in fig. 4 implements this procedure: the inputs are the predictions generated by the GRU net (fig. 2) Predictions, the indices of the entries where velocoity jumps begin PosBeforeJ and after PosAfterJ, and the real magnitude of the friction $\mathsf{\Delta F_{data}}$, the output is the evaluation of eq. 11.
Figure 4: Sub-net that implements direct effect.
Figure 5: Sub-net that implements logarithmic healing during holds.
Finally, interface strengthening during hold periods (a.k.a. healing) is recognized to happen logarithmically with respect to hold time, i.e., counting from the moment that the interface ceases to move. Consider that, given a sequence, the hold period starts at the $j$-th entry, and it lasts until $(j+N_{h})$-th entry. We can gather these terms in a separate vector $\delta\bm{\mu}_{h}=\left\{\delta\mu_{n}\right\}_{i=k}^{i=k+N_{h}}=\left\{\delta\mu_{n}\right\}_{n=1}^{n=N_{h}}$, where the sub-index $n$ has been introduced to index the values during hold straightforwardly. So the healing condition can be expressed as:
${\delta\mu_{n}-\delta\mu_{N_{h}}\over\log\left({n\over N_{h}}\right)}=\mathrm{constant}\quad\forall n=1,\ldots,N_{h}-1\,.$ (12)
Thus, the loss terms is:
$\mathsf{loss}_{4}=\sum_{n=1}^{N_{h}-2}\left\{{\delta\mu_{n+1}-\delta\mu_{N_{h}}\over\log\left({n+1\over N_{h}}\right)}-{\delta\mu_{n}-\delta\mu_{N_{h}}\over\log\left({n\over N_{h}}\right)}\right\}^{2}\,.$ (13)
This loss term is realized schematically as in fig. 5. The first input is again Predictions, coming from the main network fig. 2, the second one is $\mathsfit{N_{h}}$, which is the position of the entry where the hold ends, it is necessary to extract the predicted value $\delta\mu_{N_{h}}$. The latter is then replicated to form a vector that is subtracted to all predictions. The final input LogTerms weights the healing values as per eq. 12 and “masks” the not-healing ones since it contains zeros that multiply those entries. Once the vector has been weighted, two copies of it are made, shifted to compare successive entries and then summed and squared, yielding as output eq. 13.
The overall loss term that guides the optimization of weights and biases during the training process is:
$\mathsf{loss}=\mathsf{loss}_{1}+\alpha\mathsf{loss}_{2}+\beta\mathsf{loss}_{3}+\gamma\mathsf{loss}_{4}\,,$ (14)
$\alpha,\,\beta,\,\gamma\in\mathbb{R}$. We found $\alpha=1$, $\beta=0.1$ and $\gamma=5\cdot 10^{-3}$ to yield satisfactory results, yet probably not optimal (see discussion section).
Figure 6: Scheme of the global training network
III Training results
The training was performed in Mathematica [44], using the architecture and hyperparameters discussed in section II and 12000 random sequences corresponding to virtual protocols.
The overall training procedure is depicted in fig. 7 and the final losses are consigned in table 1. The complex loss function eq. 13 was not used at once, but an incremental approach was explored: first, the simplest loss (mean squared error, $\mathsf{ME}$) was used to initialize weights and biases. Second, the network was retrained adding the initial-condition term ($\mathsf{ME+ics}$), which further reduced the loss by 50%. This main part was again extracted and used into a larger scheme that included the direct effect loss term ($\mathsf{ME+ics+de}$). The overall precision suffered, but, as it will be shown, this deterioration of the objective function minimization does not translate into a substantial deterioration of the overall approximation; rather, it seems that this larger value reflect the struggle to capture the jumps’ magnitude. We tried re-training the network from $\mathsf{ME+ics}$ for both direct effect and healing simultaneously ($\mathsf{ME+ics+de+heal}$), to no avail. Apparently, the addition of the healing phenomenology to the loss further deflect from the optimum, but it is believed that this is really a drawback of the RSF model used to generate the synthetic data.
Figure 7: Mean training loss: $\mathsf{ME}$ is trained without initialization, $\mathsf{ME+ics}$ is trained with paramters from $\mathsf{ME}$, $\mathsf{ME+ics+de}$ and $\mathsf{ME+ics+de+heal}$ are trained with parameters from $\mathsf{ME+ics}$.
Table 1: Final loss value for different loss functions: ME refers to mean squared error between predictions and targets, ICs refers to enforcing the first value of predictions to be equal to targets’, DE does to enforces friction jumps in target data and predictions to be of the same magnitude, H refers to enforcing logarithmic healing.
ME ME+ICs ME+ICs+DE [^1] Initial parameters from ME+ICs ME+ICs+DE+H [^2] Initial parameters from ME+ICs
0.064 0.033 0.255 0.271
IV Test
Once the training is complete, we take the main net, which is our ready-to-use $\mathsf{NN}$ function that is to model the interface evolution. We test it using the 3000 sequences that the net did not see during training. For each sequence, the percent error is computed as
$\mathsf{error}(\%)=100{||\mathsf{NN}(\bm{V}/V_{ch})-\delta\bm{\hat{\mu}}/\mu_{ch}||_{2}\over||\delta\bm{\hat{\mu}}/\mu_{ch}||_{2}}\,.$ (15)
The mean test error (averaged over the 3000 sequences) is consigned in table 2 for the four different architectures discussed above. It remains fairly constant, always below 5%, the minimum being attained for the loss combining the averaged square error for the whole vector with the direct enforcement of zero change at $t=0$. Their accuracy can be directly grasped in figs. 10 and 10, which will be commented in the next section.
Finally, fig. 8 compares the error distribution across sequences for both the simplest network and the most complex one. In the latter case, the most current error is $\sim 3\%$ (19% of tests), while $\sim 2\%$ (16% of tests) in the former.
Table 2: Mean test error (%) for different loss functions (see table 1 for labels’ explanation).
ME ME+ICs ME+ICs+DE ME+ICs+DE+H
4.2 3.4 3.655 4.1
Figure 8: Verification error breakdown: percentage of test cases that yield a certain percent error.
V Discussion
As hinted by the test results presented in table 2, figs. 10 and 10 clearly showcases that the GRU-based network $\mathsf{NN}$ is able of closely reproducing the friction dynamics encapsulated in the datasets generated with a customary rate-and-state phenomenological law.
Nevertheless, there are obvious limitations. First, the inset in fig. 10 reveals that the enforcement of the initial condition prevents an initial substantial error, but it does not prevent oscillations even though the velocity remains constant. This issue is deemed of minor importance, since real interface data will certainly come with small levels of noise that the network must filter through. By the way, at this point we will have to consider also the use of regularization to prevent overfitting, a feature that was not necessary as the datasets were noiseless.
The fact that, out of the hold periods, the target (friction coefficient) changes when there is no change in the features (loading velocities) is a clear drawback, no matter how slightly the change may be.
Moreover, the absence of the state variable in the dataset brings about inaccuracies in reproducing behavior during holds, despite (or rather, because of) explicitly implementing logarithmic healing in the objective function. Logarithmic healing, eq. 8, entails that the increment of static friction coefficient $\Delta\mu_{s}$ over a timespan $\Delta t$ abides by
$\Delta\mu_{s}\propto{\Delta t\over t}\,,$ (16)
where $t$ is the total time elapsed from the arrest. Conversely, if we integrate eq. 2 and then plug it back into eq. 3, a term depending on the initial value of the state variable appears:
$\Delta\mu_{s}\propto{\Delta t\over\theta(t=0)+t}\,,$ (17)
which can interfere with the scaling, unless $\theta(t=0)\ll t$. The latter condition may be met in real experiments, in which the holds tend to last for much longer than the characteristic age of the interface, but it is certainly not the case here as the holds are always “short”, i.e., of duration similar magnitude to the magnitude of the state variable. This seems to be the most obvious drawback of not considering explicitly the interface state in the data-driven framework. In any case, it does not appear as insurmountable, since (as justified above)
Yet another aspect to acknowledge is that the magnitude of direct effect peaks seems to be affected by the holds. See that in fig. 10 there is a relatively short hold period preceded and ended by jumps which are well captured by the nets; conversely, the jumps’ magnitude are significantly underpredicted in fig. 10, which features a longer hold, which particularly the net that accounts explicitly for healing struggles to capture.
Some of these issues could be remedied drawing inspiration from PINNs [29] again, by taking advantage of the automatic differentiation [47] to automatically compute gradients of the friction coefficient (output) as a function of the velocities (input) so that they can be directly monitored during training. Not only that, it is logical to expect that automatic differentiation can also enable more elegant approaches to implementing both direct effect and healing in the loss function.
Figure 9: Test example #1: light left axis for loading velocity, dark right axis for friction coefficient evolution. Emphasis on initial conditions. The dashed background line represents the velocity protocol. Protocol with hold playing a secondary role.
Figure 9: Test example #1: light left axis for loading velocity, dark right axis for friction coefficient evolution. Emphasis on initial conditions. The dashed background line represents the velocity protocol. Protocol with hold playing a secondary role.
Figure 10: Test example #2: light left axis for loading velocity, dark right axis for friction coefficient evolution. Emphasis on shortcomings in healing and direct effect. The dashed background line represents the velocity protocol. Protocol with relevant hold.
VI Final Remarks
This study has demonstrated that recurrent neural networks based on Gated Recurrent Units (GRUs) possess the capability to accurately model slide-hold protocols in the context of rate-and-state friction laws, provided that the networks are supplied with adequate data and the loss function is tuned accordingly. Our findings affirm the potential of recurrent neural networks as a powerful tool for understanding complex frictional behavior, showcasing their ability to learn from and predict the dynamics of friction with high fidelity. This success highlights the ability of deep learning approaches to replace customary phenomenological laws for dynamic friction, additionally avoiding the need to define internal variables.
However, our investigation also identified limitations, particularly concerning the handling of logarithmic healing during short hold periods. This limitation suggests that future models may incorporate mechanisms to better account for short holds, thereby enhancing their accuracy and applicability to a broader range of slip events. Additionally, we observed a potential conflict between optimizing for healing and for the direct effect within the training process, indicating an area for further refinement in how these fundamental aspects of frictional behavior are integrated into the learning scheme. And let alone that no noise has been included in the training datasets yet, and only protocols made of sudden velocity jumps changes were considered.
This document contains but an exploratory approach to modeling frictional interface via deep learning. Looking forward, there are several promising avenues for enhancing the predictive power and generalizability of recurrent neural networks for dynamic friction. Improvement could be sought through the exploration of alternative RNN architectures and more nuanced physics-informed loss functions (for instance, to account for friction manifesting indirectly in the dynamics of spring-block systems [2]). Further parametric studies in terms of network hyperparameters are also warranted, in particular in terms of volumes of available data: unlike cheap synthetic data, experimental data remains relatively scarce; hence it is worth exploring how new techniques [48, 49], aimed at improving data economy in the context of constitutive modeling with neural networks,
As mentioned above, incorporating spring-block dynamics into the model could pave the way for learning from real experimental data, thus circumventing the gap between theoretical models and practical observations. Combined with the use of automatic differentiation [47] (not explored at this point), could offer a completely new pathway to take advantage of experimental data to understand friction, replacing the use of experiments to simply fit parameters by producing training data that is the infused in a network that can be used later for predictions.
Moreover, the exploration of advanced machine learning features such as transformers [50] and Long Short-Term Memory (LSTM) networks [51] could offer additional improvements. These technologies, known for their ability to capture long-distance dependencies within data, might provide the key to overcoming the current limitations. The promise of deep learning does not stop there: one can foresee that an approach based on neural operators [52] may help also with the problem of generalizing friction from lab conditions to crustal ones, the network operator may be brought to bear in this case.
The ultimate objective is to transition from conventional phenomenological models of dynamic friction to more faithful representations directly learned from the same data traditionally used for parameter fitting. This shift should not entail a departure from our pursuit of a fundamental understanding of friction dynamics. On the contrary, it necessitates a more integrated approach where firm physical principles are embedded within the deep learning framework. The adoption of these advanced models should be driven by their proven efficacy in capturing interface behaviors, thereby enhancing their utility in larger-scale simulations where interface response plays a role. This strategy aims not to replace the foundational theories of friction but to complement them with data-driven insights, ensuring a more comprehensive and precise depiction of frictional phenomena to be included in numerical analyses.
Acknowledgements
The author acknowledges financial support from the Swiss National Science Foundation, via Ambizione Grant PZ00P2_216341 “Data-Driven Computational Friction”. Useful conversations (regarding dynamic friction, rate-and-state friction and/or laboratory experiments) with Dr. Gabriel Meyer, Dr. Simon Guérin-Marthe and Ms. Roxane Ferry are gratefully acknowledged. An EPFL bachelor project conducted by Mr. Antoine Binggeli provided Tensorflow code that inspired the use of GRUs to this particular problem.
[…]
Data availability
The notebooks used to generate both data and results in this paper are available via correspondence with the author (joaquin.garciasuarez@epfl.ch).
Appendix A Alternative loss for direct effect
An alternative loss function to enforce direct effect, inspired by the one used for healing, is also presented. It has not been implemented in the previous analyses, but it is shown for completeness’ sake.
The magnitude of the jump in friction coefficient must be proportional to the logarithm of the ratio of final and initial velocities. Mathematically, we express that as
$\mu(V_{\text{a}})-\mu(V_{\text{b}})\propto$ $\log\left({V_{a}\over V_{b}}\right)$
$\to\delta\mu_{\text{a}}-\delta\mu_{\text{b}}\propto$ $\log\left({V_{a}\over V_{b}}\right)\,,$ (18)
where the sub-indices $a$ and $b$ refer to after and before the velocity jump. This is a physical constraint that does not have to be directly referred to the target values, but it is expressed as a condition that the predictions have to abide by. Thus, for all jumps during a virtual experiment, we should have
$\left[{\delta\mu_{\text{a}}-\delta\mu_{\text{b}}\over\log\left({V_{a}\over V_{b}}\right)}\right]_{J_{1}}=\ldots=\left[{\delta\mu_{\text{a}}-\delta\mu_{\text{b}}\over\log\left({V_{a}\over V_{b}}\right)}\right]_{J_{N_{J}}}\,,$ (19)
where $N_{J}$ refers to the total number of velocity jumps in any given velocity protocol, so $J_{1}$ means “first jump” and $J_{N_{J}}$ “last jump”. Hence, the loss term that promotes the enforcement of the direct effect, expressed in this way, is:
$\mathsf{loss}_{3}=\sum_{i=1}^{N_{J}-1}\left\{\left[{\delta\mu_{\text{a}}-\delta\mu_{\text{b}}\over\log\left({V_{a}\over V_{b}}\right)}\right]_{J_{i+1}}-\left[{\delta\mu_{\text{a}}-\delta\mu_{\text{b}}\over\log\left({V_{a}\over V_{b}}\right)}\right]_{J_{i}}\right\}^{2}\,,$ (20)
in other words, the magnitude of the friction coefficient change over any velocity jump must be the same across jumps.